A Bibliography of Papers in Lecture Notes in Computer Science (2010)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 April 2024 Version 1.06

Title word cross-reference

L [10].

1.0 [45]. 10th [66]. 11th [63]. 12th [65]. 16th [58].

2 [24]. 2.5 [36]. 29th [61].

30th [67].

5th [59].

8.8/11.2 [62].

9th [62, 64].

Accurate [22]. advanced [62]. Advances [58, 61, 67]. Africa [60]. Africacrypt [60]. Algebra [15, 12, 37, 41, 24]. Algebraic [20]. Algorithm [38, 34, 39]. Algorithms [40]. Allocation [16]. Alternative [4]. America [57]. Analysis [35]. annual [61, 67]. application [58, 62]. applications [61]. April [62]. Arbitrary [38]. Arithmetic [19, 54, 22]. Asiacrypt [58]. August [57, 67]. Automated [41, 30].

Barbara [67]. Based [23, 33]. Bases [42, 10, 17]. Beijing [59]. Beyond [15]. Binomial [38]. Bivariate [21]. Black [11]. Black-Box [11]. Blocks [16]. Box [11]. Branch [34]. Building [27, 16].

C [15, 53, 56, 48]. C-XSC [56]. CA [66, 67].

CANS [64]. card [62]. CARDIS [62]. Certified [19, 52]. CGAL [20]. Characteristic [38]. China [59]. Class [53]. Classical [11]. CoCoALib [15]. Codes [5]. Coherent [14]. Cohomology [9]. Communicating [56]. Commutative [15]. Compilation [16]. Complex [53]. Computation [14, 24]. Computational [12, 1, 35]. Computations [15, 20, 52, 23]. Computer [37]. Computing [10, 17, 19, 3, 13, 39, 40, 22, 8]. Conference [62, 57, 58, 59, 60, 61, 63, 64, 65, 66, 67]. Configurations [14]. Confluence [41]. Congress [69]. Conjecture [28]. Constraints [46]. Construction [23, 32]. Constructive [11]. Controlled [19]. Core [24]. CoStLy [53]. Create [4]. Crypto [67]. cryptographer [66]. cryptographic [61]. Cryptography [68]. cryptology [57, 58, 59, 60, 61, 63, 65, 66, 67, 64]. CT [66]. CT-RSA [66].

DAG [23]. Dags [23]. Darmstadt [68]. Data [35]. DDMF [7]. December [58, 59, 63, 64, 65]. Decision [23]. Decomposition [40]. Deferring [23]. Definite [39]. Deformations [31]. Design [24]. Desingularization [38]. Development [5]. Dictionary [7]. Differential [39, 41]. Discrete [1, 35]. Distance [34]. Do [27]. Dynamic [7].

Efficient [55, 3]. Engine [50]. Enumerating [44]. Environment [5]. Equations [39]. Euler [26]. Eurocrypt [61]. Evaluation [55, 30]. Exact [20, 23, 24]. Exploiting [2, 48]. Expression [23]. Expressions [56].

Fans [47]. Fast [18]. FGb [17]. Filters [33]. Finitely [10]. first [57]. Fixed [19]. Fixed-Precision [19]. Floating [22]. Floating-Point [22]. Floats [23]. Formal [29]. Formula [26]. Founding [16].

Framework [35]. Francisco [66]. Free [4]. Function [13]. Functional [56]. Functions [7, 6, 53].

Galois [44]. GAP [9, 12, 14]. Generating [51]. Geometric [19, 20, 3]. Geometry [1, 37, 34, 32, 24]. Germany [62, 68]. GNU [8]. Gröbner [42, 10, 17]. Groups [11, 10, 13].

Harmonic [32]. High [12].

High-Performance [12]. HOL [25].

Hyderabad [63].

ICISC [65]. ICMS [69]. Ideals [40]. IFIP [62]. Image [33]. Implementation [42, 34]. Improvement [13]. Inclusion [53]. India [63]. Indocrypt [63]. Infinite [41]. information [57, 58, 59, 65]. Inhomogeneous [39]. INSCRYPT [59]. Integer [49]. Integrals [39]. Integro [41]. Integro-Differential [41]. International [69, 62, 68, 57, 58, 59, 60, 61, 63, 64, 65]. Interval [54]. Introducing [25]. Introduction [36, 18]. isl [49].

Japan [69]. jeep [34]. June [61].

Kōbe [69]. Kepler [28]. Korea [65]. Kuala [64].

Large [2, 52, 55]. Latin [57]. Latincrypt [57]. Lattice [51]. Libraries [54]. Library [15, 27, 17, 18, 53, 30, 49, 24, 33]. Like [27]. LinBox [16]. Linear [28]. Local [31]. Lumpur [64].

Magma [43, 4]. Malaysia [64]. Maple [4]. March [66]. Mathemagix [52]. Mathematica [4, 56]. Mathematical [27, 7, 4, 69, 30, 33, 69]. MATLAB [4]. May [60, 61, 68]. MD [34]. MD-jeep [34]. Mechanized [27]. Membership [11]. Meshes [31]. Methods [37]. Mexico [57].

Mizar [26, 30]. Model [49]. Monaco [61]. Morse [35]. MPFR [8].

network [64]. Normaliz [36]. Normalizers [13]. Number [43, 18]. Numeric [52, 24]. Numerical [5]. NZMATH [45].

OpenGL [33]. OpenGL-Based [33]. Operadic [42]. Optimization [50]. Orbifolds [9].

Package [14]. papers [59, 65]. Parallel [16]. Parametrized [41]. Passau [62]. Performance [12]. Permutation [13]. Perturbation [19]. Pick [29]. Planar [31]. Point [22]. Polycyclic [10]. Polyhedral [47, 48, 49]. Polyhedron [26]. Polynomial [40]. Polynomials [55, 21]. Polytopes [51]. Post [68]. Post-Quantum [68]. PQCrypto [68]. Precision [19]. Prescribed [32]. Presented [10]. Primary [40]. Problems [2, 34]. Proceedings [58, 61, 63, 64, 66, 67, 68, 57, 60, 69, 62]. Programming [2, 52]. Programs [28]. Progress [57, 60, 63]. Project [4]. Proof [29, 41]. Proofs [27]. Proving [30]. Prune [34]. Puebla [57].

Quad [31]. Quad-Meshes [31]. Quadratic [46]. Quantum [68]. Quotients [10].

Redundant [46]. Reformulation [50]. Reformulation-Optimization [50]. Reliable [3, 22, 8]. Removing [46]. Representations [44]. research [62]. revised [59, 65]. Rewrite [41]. RSA [66].

Sage [44, 4]. San [66]. Santa [67]. Scale [2, 52]. Scope [16]. security [57, 58, 59, 64, 65]. selected [59, 65]. Semidefinite [2]. Seoul [65]. Separate [16]. September [69]. Set [49]. Simplicial [9]. Singapore [58]. Smart [62]. Smooth [9, 51]. Software [6, 4, 50, 69]. Sollya [5].

Solving [21]. South [60]. Sparsity [2]. Special [6]. Speeds [23]. Speeds-Up [23]. Standardized [54]. Stellenbosch [60]. Storing [23]. Structured [2]. Sums [23]. Surfaces [32]. Symbolic [52]. Symmetric [47]. Symmetries [48]. System [41]. Systems [21].

TADD [35]. techniques [61]. Testing [11]. Theorem [29, 30]. Theory [43, 18, 35, 58, 61]. Third [68, 60, 69]. Tool [43]. Tools [48]. Topics [66]. track [66]. Traversing [47]. Tropical [37].

Unifying [4]. USA [66, 67]. Used [54].

Using [35].

Validated [6]. Varieties [38]. via [10].

Viable [4].

WG [62]. Workshop [68].

XSC [56].

Zero [25].

References

Hales:2010:CDG

[1] Thomas C. Hales. Computational discrete geometry. Lecture Notes in Computer Science, 6327:1–3, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                            link.springer.com/content/pdf/bfm:
                                            978-3-642-15582-6/1.pdf;   http:/
                                            /link.springer.com/content/pdf/10.
                                            1007/978-3-642-15582-6_1.pdf.

Kojima:2010:ESS

[2] Masakazu Kojima. Exploiting structured sparsity in large scale semidefinite programming problems. Lecture Notes in Computer Science, 6327:4–9, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_2.pdf.

Mehlhorn:2010:REG

[3] Kurt Mehlhorn. Reliable and efficient geometric computing. Lecture Notes in Computer Science, 6327:10– 11, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_3.pdf.

Erocal:2010:SPU

[4] Burçin Eröcal and William Stein. The Sage Project: Unifying free mathematical software to create a viable alternative to Magma, Maple, Mathematica and MAT-LAB. Lecture Notes in Computer Science, 6327:12–27, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_4.pdf.

Chevillard:2010:SED

[5] Sylvain Chevillard, Mioara Joldeş, and Christoph Lauter. Sollya: an environment for the development of numerical codes. Lecture Notes in Computer Science, 6327:28–31, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-15582-6_5.pdf.

Cuyt:2010:VSF

[6] Annie Cuyt, Franky Backeljauw, Stefan Becuwe, and Joris Van Deun. Validated special functions software.

Lecture Notes in Computer Science, 6327:32–34, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/978-3-642-
                                        15582-6_6.pdf.

Benoit:2010:DDM

[7] Alexandre Benoit, Frédéric Chyzak, Alexis Darrasse, Stefan Gerhold, and Marc Mezzarobba …. The Dynamic Dictionary of Mathematical Functions (DDMF). Lecture Notes in Computer Science, 6327:35–41, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        978-3-642-15582-6_7.pdf.

Zimmermann:2010:RCG

[8] Paul Zimmermann. Reliable computing with GNU MPFR. Lecture Notes in Computer Science, 6327:42–45, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-15582-6_8.pdf.

Barakat:2010:SCS

[9] Mohamed Barakat and Simon Görtzen. Simplicial cohomology of smooth orbifolds in GAP. Lecture Notes in Computer Science, 6327:46–49, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-15582-6_9.pdf.

Eick:2010:CPQ

[10] Bettina Eick and Max Horn. Computing polycyclic quotients of finitely (L- )Presented groups via Gröbner bases. Lecture Notes in Computer Science, 6327:50–53, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_10.pdf.

Ambrose:2010:CMT

[11] Sophie Ambrose, Scott H. Murray, Cheryl E. Praeger, and Csaba Schneider. Constructive membership testing in black-box classical groups. Lecture Notes in Computer Science, 6327:54–57, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_11.
pdf.

Behrends:2010:THP

[12] Reimer Behrends, Alexander Konovalov, Steve Linton, Frank Lübeck, and Max Neunhöeffer. Towards high-performance computational algebra with GAP. Lecture Notes in Computer Science, 6327:58– 61, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_12.
pdf.

Miyamoto:2010:IFC

[13] Izumi Miyamoto. An improvement of a function computing normalizers for permutation groups. Lecture Notes in Computer Science, 6327:62–68, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_13.pdf.

Pasechnik:2010:GPC

[14] Dmitrii V. Pasechnik and Keshav Kini.

A GAP package for computation with coherent configurations. Lecture Notes in Computer Science, 6327:69–72, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-15582-6_14.
                                        pdf.

Abbott:2010:CCL

[15] John Abbott and Anna M. Bigatti. Co-CoALib: a C++ library for computations in commutative algebra … and beyond. Lecture Notes in Computer Science, 6327:73–76, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        978-3-642-15582-6_15.pdf.

Dumas:2010:LFS

[16] Jean-Guillaume Dumas, Thierry Gautier, Clément Pernet, and B. David Saunders. LinBox founding scope allocation, parallel building blocks, and separate compilation. Lecture Notes in Computer Science, 6327:77–83, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/978-3-642-
                                        15582-6_16.pdf.

Faugere:2010:FLC

[17] Jean-Charles Faugère. FGb: a library for computing Gröbner bases. Lecture Notes in Computer Science, 6327:84–87, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-15582-6_17.
                                        pdf.

Hart:2010:FLN

[18] William B. Hart. Fast library for number theory: an introduction. Lecture Notes in Computer Science, 6327:88–91, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_18.
pdf.

Halperin:2010:CPC

[19] Dan Halperin. Controlled perturbation for certified geometric computing with fixed-precision arithmetic. Lecture Notes in Computer Science, 6327:92–95, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_19.
pdf.

Karavelas:2010:EGA

[20] Menelaos I. Karavelas. Exact geometric and algebraic computations in CGAL. Lecture Notes in Computer Science, 6327:96–99, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_20.pdf.

Rouillier:2010:SSB

[21] Fabrice Rouillier. On solving systems of bivariate polynomials. Lecture Notes in Computer Science, 6327:100–104, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_21.
pdf.

Rump:2010:ARC

[22] Siegfried M. Rump. Accurate and reliable computing in floating-point arithmetic. Lecture Notes in Computer Science, 6327:105–108, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-15582-6_22.pdf.

Morig:2010:DDC

[23] Marc Mörig. Deferring DAG construction by storing sums of floats speeds-up exact decision computations based on expression dags. Lecture Notes in Computer Science, 6327:109–120, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_23.pdf.

Yu:2010:DCL

[24] Jihun Yu, Chee Yap, Zilin Du, Sylvain Pion, and Hervé Brönnimann. The design of Core 2: a library for exact numeric computation in geometry and algebra. Lecture Notes in Computer Science, 6327:121–141, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_24.pdf.

Adams:2010:IHZ

[25] Mark Adams. Introducing HOL Zero. Lecture Notes in Computer Science, 6327:142–143, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_25.pdf.

Alama:2010:EPF

[26] Jesse Alama. Euler’s polyhedron formula in Mizar. Lecture Notes in Computer Science, 6327:144–147, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_26.pdf.

Arthan:2010:BLM

[27] R. D. Arthan. Building a library of mechanized mathematical proofs: Why do it? and what is it like to do? Lecture Notes in Computer Science, 6327:148, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_27.pdf.

Hales:2010:LPK

[28] Thomas C. Hales. Linear programs for the Kepler conjecture. Lecture Notes in Computer Science, 6327:149–151, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_28.
pdf.

Harrison:2010:FPP

[29] John Harrison. A formal proof of Pick’s theorem. Lecture Notes in Computer Science, 6327:152–154, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-15582-6_29.pdf.

Urban:2010:EAT

[30] Josef Urban, Krystof Hoder, and Andrei Voronkov. Evaluation of automated theorem proving on the Mizar mathematical library. Lecture Notes in Computer Science, 6327:155–166, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_30.pdf.

Hoffmann:2010:LDP

[31] Tim Hoffmann. On local deformations of planar quad-meshes. Lecture Notes in Computer Science, 6327:167–169, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_31.
pdf.

Weber:2010:CHS

[32] Matthias Weber. Construction of harmonic surfaces with prescribed geometry. Lecture Notes in Computer Science, 6327:170–173, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_32.pdf.

vonGagern:2010:LOB

[33] Martin von Gagern and Christian Mercat. A library of OpenGL-based mathematical image filters. Lecture Notes in Computer Science, 6327:174–185, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_33.pdf.

Mucherino:2010:MJI

[34] Antonio Mucherino, Leo Liberti, and Carlile Lavor. MD-jeep: an implementation of a branch and prune algorithm for distance geometry problems. Lecture Notes in Computer Science, 6327:186– 197, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_34.
pdf.

Reininghaus:2010:TCF

[35] Jan Reininghaus, David Günther, Ingrid Hotz, Steffen Prohaska, and Hans-Christian Hege. TADD: a computational framework for data analysis using discrete Morse theory. Lecture Notes in Computer Science, 6327:198–208, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_35.pdf.

Bruns:2010:IN

[36] Winfried Bruns, Bogdan Ichim, and Christof Söger. Introduction to Normaliz 2.5. Lecture Notes in Computer Science, 6327:209–212, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_36.pdf.

Markwig:2010:CAM

[37] Thomas Markwig. Computer algebra methods in tropical geometry. Lecture Notes in Computer Science, 6327:213– 216, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_37.
pdf.

Blanco:2010:NDA

[38] Rocío Blanco. A new desingularization algorithm for binomial varieties in arbitrary characteristic. Lecture Notes in Computer Science, 6327:217–220, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_38.pdf.

Nakayama:2010:ACI

[39] Hiromasa Nakayama and Kenta Nishiyama. An algorithm of computing inhomogeneous differential equations for definite integrals. Lecture Notes in Computer Science, 6327:221–232, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-15582-6_39.pdf.

Noro:2010:NAC

[40] Masayuki Noro. New algorithms for computing primary decomposition of polynomial ideals. Lecture Notes in Computer Science, 6327:233–244, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_40.pdf.

Tec:2010:ACP

[41] Loredana Tec, Georg Regensburger, Markus Rosenkranz, and Bruno Buchberger. An automated confluence proof for an infinite rewrite system parametrized over an integro-differential algebra. Lecture Notes in Computer Science, 6327:245–248, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-15582-6_41.pdf.

Dotsenko:2010:OGB

[42] Vladimir Dotsenko and Mikael Vejdemo-Johansson. Operadic Gröbner bases: an implementation. Lecture Notes in Computer Science, 6327:249–252, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_42.pdf.

Cannon:2010:MTN

[43] John Cannon, Steve Donnelly, Claus Fieker, and Mark Watkins. Magma — a tool for number theory. Lecture Notes in Computer Science, 6327:253– 255, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_43.
pdf.

Citro:2010:EGR

[44] Craig Citro and Alexandru Ghitza. Enumerating Galois representations in Sage. Lecture Notes in Computer Science, 6327:256–259, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_44.pdf.

Tanaka:2010:N

[45] Satoru Tanaka, Naoki Ogura, Ken Nakamula, Tetsushi Matsui, and Shigenori Uchiyama. NZMATH 1.0. Lecture Notes in Computer Science, 6327:260– 269, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_45.
pdf.

Adjiashvili:2010:RRQ

[46] David Adjiashvili, Michel Baes, and Philipp Rostalski. Removing redundant quadratic constraints. Lecture Notes in Computer Science, 6327:270–281, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-15582-6_46.
                                        pdf.

Jensen:2010:TSP

[47] Anders Nedergaard Jensen. Traversing symmetric polyhedral fans. Lecture Notes in Computer Science, 6327:282– 294, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-15582-6_47.
                                        pdf.

Rehn:2010:CTE

[48] Thomas Rehn and Achill Schürmann. C++ tools for exploiting polyhedral symmetries. Lecture Notes in Computer Science, 6327:295–298, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        978-3-642-15582-6_48.pdf.

Verdoolaege:2010:IIS

[49] Sven Verdoolaege. isl: an integer set library for the polyhedral model. Lecture Notes in Computer Science, 6327:299– 302, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-15582-6_49.
                                        pdf.

Liberti:2010:ROS

[50] Leo Liberti, Sonia Cafieri, and David Savourey. The reformulation-optimization software engine. Lecture Notes in Computer Science, 6327:303–314, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_50.pdf.

Haase:2010:GSL

[51] Christian Haase, Benjamin Lorenz, and Andreas Paffenholz. Generating smooth lattice polytopes. Lecture Notes in Computer Science, 6327:315–328, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_51.pdf.

Lecerf:2010:MTL

[52] Grégoire Lecerf. Mathemagix: Towards large scale programming for symbolic and certified numeric computations. Lecture Notes in Computer Science, 6327:329– 332, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_52.
pdf.

Neher:2010:CIF

[53] Markus Neher. Complex inclusion functions in the CoStLy C++ class library. Lecture Notes in Computer Science, 6327:333–336, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
15582-6_53.pdf.

Revol:2010:SIA

[54] Nathalie Revol. Standardized interval arithmetic and interval arithmetic used in libraries. Lecture Notes in Computer Science, 6327:337–341, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_54.pdf.

Leiserson:2010:EEL

[55] Charles E. Leiserson, Liyun Li, Marc Moreno Maza, and Yuzhen Xie. Efficient evaluation of large polynomials. Lecture Notes in Computer Science, 6327:342– 353, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_55.
pdf.

Popova:2010:CFE

[56] Evgenija D. Popova and Walter Krämer. Communicating functional expressions from Mathematica to C-XSC. Lecture Notes in Computer Science, 6327:354– 365, 2010. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-15582-6_56.
pdf.

Abdalla:2010:PCL

[57] Michel Abdalla and Paulo S. L. M. Barreto, editors. Progress in cryptology — Latincrypt 2010: first international conference on cryptology and information security in Latin America, Puebla, Mexico, August 8–11, 2010, proceedings, volume 6212 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-14711-9 (softcover). LCCN ????

Abe:2010:ACA

[58] Masayuki Abe, editor. Advances in cryptology — Asiacrypt 2010: 16th international conference on the theory and application of cryptology and information security, Singapore, December 5–9, 2010. Proceedings, volume 6477 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-17372-1 (softcover). LCCN ????

Bao:2010:ISC

[59] Feng Bao, Moti Yung, and Dongdai Lin, editors. Information security and cryptology: 5th international conference, IN-SCRYPT 2009, Beijing, China, December 12–15, 2009. revised selected papers, volume 6151 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-16341-6 (softcover). LCCN ????

Bernstein:2010:PCA

[60] Daniel J. Bernstein and Tanja Lange, editors. Progress in cryptology — Africacrypt 2010: third international conference on cryptology in Africa, Stellenbosch, South Africa, May 3–6, 2010. proceedings, volume 6055 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-12677-4 (softcover). LCCN ????

Gilbert:2010:ACE

[61] Henri Gilbert, editor. Advances in cryptology — Eurocrypt 2010: 29th annual international conference on the theory and applications of cryptographic techniques, Monaco, May 30–June 3, 2010. Proceedings, volume 6110 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-13189-1 (softcover). LCCN ????

Gollmann:2010:SCR

[62] Dieter Gollmann, Jean-Louis Lanet, and Julien Iguchi-Cartigny, editors. Smart card research and advanced application: 9th IFIP WG 8.8/11.2 International Conference, CARDIS 2010, Passau, Germany, April 14–16, 2010: proceedings, volume 6035 of Lecture Notes in Computer Science. Springer, Berlin, Germany, 2010. ISBN 3-642-12509-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.S62 C36 2010.

Gong:2010:PCI

[63] Guang Gong and Kishan Chand Gupta, editors. Progress in cryptology — Indocrypt 2010: 11th international conference on cryptology in India, Hyderabad, India, December 12–15, 2010. Proceedings, volume 6498 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-17400-0 (softcover). LCCN ????

Heng:2010:CNS

[64] Swee-Huay Heng, Rebecca N. Wright, and Bok-Min Goi, editors. Cryptology and network security: 9th international conference, CANS 2010, Kuala Lumpur, Malaysia, December 12–14, 2010. Proceedings, volume 6467 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-17618-6 (softcover). LCCN ????

Lee:2010:ISC

[65] Donghoon Lee and Seokhie Hong, editors. Information, security and cryptology – ICISC 2009: 12th international conference, Seoul, Korea, December 2– 4, 2009, revised selected papers, volume 5984 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-14422-5 (softcover). LCCN ????

Pieprzyk:2010:TCC

[66] Josef Pieprzyk, editor. Topics in cryptology — CT-RSA 2010: the 10th cryptographers’ track at the RSA conference 2010, San Francisco, CA, USA, March 1– 5, 2010. Proceedings, volume 5985 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-11924-7 (softcover). LCCN ????

Rabin:2010:ACC

[67] Tal Rabin, editor. Advances in cryptology — Crypto 2010: 30th annual cryptology conference, Santa Barbara, CA, USA, August 15–19, 2010. Proceedings, volume 6223 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-14622-8 (softcover). LCCN ????

Sendrier:2010:PQC

[68] Nicolas Sendrier, editor. Post-Quantum Cryptography: Third International Workshop, PQCrypto 2010, Darmstadt, Germany, May 25–28, 2010. Proceedings, volume 6061 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-12928-5. LCCN QA76.9.A25 2010.

Fukuda:2010:MSI

[69] Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama, editors. Mathematical software — ICMS 2010: third International Congress on Mathematical Software, Kōbe, Japan, September 13–17, 2010: proceedings, volume 6327 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 3-642-15581-2 (paperback), 3-642-15582-0 (e-book). LCCN QA76.95 .I5654

                                        2010. URL http://www.springerlink.
                                        com/content/978-3-642-15581-9.